OpenStudy (ttop0816):

The range of the function y = secx - 2 is all reals except -3

OpenStudy (tkhunny):

y = cos(x) [-1,1] y = sec(x) \((−\infty,−1]∪[1,\infty)\) y = sec(x) - 2 You tell me!

OpenStudy (ttop0816):

@tkhunny is that a no....?

OpenStudy (ttop0816):

):

OpenStudy (tkhunny):

I didn't say that. Why do you doubt. Did you get it or not?

OpenStudy (anonymous):

What does sec equate to?

OpenStudy (ttop0816):

@nelsonjedi 1/cos

OpenStudy (anonymous):

Correct..now what is the range for cos?

OpenStudy (tkhunny):

It doesn't "equal to" anything. It is a function requiring an argument. Generally, we see this definition: \(\sec(x) = \dfrac{1}{\cos(x)}\)

OpenStudy (tkhunny):

Please don't EVER write 'sec' or 'cos' in a formula. Just don't do it. You THINK it's a convenient shorthand. It's really just confusing.

OpenStudy (ttop0816):

@tkhunny oh.... okay! then sec(x)=1/cos(x) @nelsonjedi (-1,1)!

OpenStudy (tkhunny):

How do y = sec(x) and y = sec(x) - 2 compare? The Domain of y = sec(x) is everything Real except (-1,1) Move that down two. You had it right. I'm just trying to get you to prove it and be SURE!

OpenStudy (ttop0816):

@tkhunny oh! hahahaha then my answer if correct?? :D

OpenStudy (tkhunny):

Yes it is. You DO need to understand why so that you can go confidently forward.